Solve the exponential equation by expressing each side as a power of the same base and then equating exponents.
(x-8)/(3^8) = sqrt ( 3)
The solution set is { }

Answer by Theo(13342)

your equation is:

(x-8) / 3^8 = sqrt(3)

since sqrt(3) = 3^(1/2), your equation becomes:

(x-8) / 3^8 = 3^(1/2)

(x-8) = 3^y if and only if log3(x-8) = y

this means that (x-8) = 3^(log3(x-8))

your equation becomes:

3^(log3(x-8)) / 3^8 = 3^(1/2)

since 3^(log3(x-8)) / 3^8 is equivalent to 3^(log3(x-8) - 8), your equation becomes:

3^(log3(x-8)-8) = 3^(1/2)

this is true if and only if log3(x-8)-8) = 1/2

add 8 to both sides to get log3(x-8) = 8 + 1/2

simplify to get log3(x-8) = 17/2

this is true if and only if 3^(17/2) = x-8

add 8 to both sides of this equation to get x = 8 + 3^(17/2)

to confirm the solution is good, go back to the original equation and replace x with (8 + 3^(17/2)

the original equation is (x-8) / 3^8 = sqrt(3)

after you replace x, the equation becomes:

(8 + 3^(17/2) - 8) / 3^8 = 3^(1/2

the 8 - 8 cancels out and you are left with:

3^(17/2) / 3^8 = 3^(1/2)

since x^a / x^b = x^(a-b), this equation becomes:

3^((17/2)-8) = 3^(1/2)

make common denominators of the fractions to get:

3^((17/2)-(16/2)) = 3^(1/2)

combine like terms to get:

3^(1/2) = 3^(1/2)

this confirms the solution is correct.

the solution is x = 8 + 3^(17/2)

this confirms the solution is correct.

another way you could have solved it is as follows:

start with:
(x-8)/3^8 = 3^(1/2)

multiply both sides of this equation by 3^8 to get:

x-8 = 3^(1/2) * 3^8)

this becomes x-8 = 3^(17/2)

add 8 to both sides of this eqution to get:

x = 3^(17/2) + 8

this way is much quicker but the problem stated that everything had to be put in base 3 to an exponent form and so there was more to do in order to accomplish that.

either way you get the same answer which is always a good thing.